# Bridging or Breaking: Impact of Intergroup Interactions on Religious Polarization

### 20 Feb 2024 v2

@@ Abstract.
Social media, Polarization, Intergroup Interaction, Religion, Heterogeneous Treatment Effects
copyright: nonejournalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645675isbn: 979-8-4007-0171-9/24/05conference: the ACM Web Conference 2024; May 13–17, 2024; Singaporeccs: Information systems Social networksccs: Social and professional topics Religious orientationccs: Applied computing Economicsccs: Applied computing Sociology
@@ 1. Introduction
Polarization between identity groups is an important driver of social unrest and may adversely affect a nation’s economic growth and responses to crises (Esteban and Schneider, 2008; Egorov et al., 2021). However, it is less clear how polarization evolves during crises. On the one hand, collective suffering may foster within-group solidarity (Bauer et al., 2016). On the other hand, it may lead to the attribution of blame on the “outside” group or increase between-group competition for limited resources (Homer-Dixon, 2010). These behaviors might be accentuated among people who have a tendency to interact only within their groups and hence might have a restricted information environment (Halberstam and Knight, 2016). Polarization has been widely studied in the social-media context as well. In particular, social media platforms such as Twitter have become more polarized over time (Garimella and Weber, 2017). Consistent with this, political polarization decreased for users who deactivated their Facebook accounts before the US elections (Allcott et al., 2020). The depolarization, however, depends on an individual’s background and this pattern might reverse for individuals having homogeneous offline social networks (Asimovic et al., 2021). Interactions within social media platforms might have varying impacts on polarization. In an unfavourable environment, exposure to outgroup viewpoints may lead to stereotype formation and increase polarization. For example, Bail et al. (2018) conduct a field experiment and find that Republicans (Democrats) who were offered financial incentives to follow a liberal (conservative) Twitter bot became more conservative (liberal).
@@ 3.1. Inferring Religion
We drop verified users from our data to remove influential individuals/organizations. To reduce possible biases in the polarization measure due to erroneous classification of religion, we remove non-personal names. We discuss the details of name cleaning steps in Appendix S2. Our final data comprises 692,559 unique users with 3,072,503 tweets. [^6] To check for the possibility that our results might be influenced by bots, we use the recent lists TwiBot-20 (Feng et al., 2021) and TwiBot-22 (Feng et al., 2022). We find that less than 0.5% of users in our data are listed as bots and contribute to 0.63% of the tweets. The model from (Chaturvedi and Chaturvedi, 2024) is trained on names obtained from a nationally representative sample, while Twitter usernames tend to be noisy. The name classification exercise also depends on the distribution of names in the specific domain the algorithm was trained on. Therefore, we expect a domain shift when applying the model to our data. To address this, we manually annotate a subset of approximately one thousand names as Muslim or non-Muslim. Since Muslim and non-Muslim names are linguistically distinct, prior literature has also relied on manually annotating religion as Muslim and non-Muslim from names in the South Asian context. In Appendix A, we discuss how we select the sample of names for manual annotation and choose a more suitable classification threshold. Chaturvedi and Chaturvedi (2024) report an $F_{1}$ score of 95% on their test set which reduces to around 85% when only one name part is available (i.e., only the first or last name)—as is often the case on Twitter; we find comparable values (sensitivity 84% and specificity 86.5%). The misclassifications lead to measurement error, potentially underestimating polarization; though qualitatively we expect the temporal patterns to remain the same. Therefore, our estimates are likely to be conservative estimates. Similarly, since we focus on changes in GCS over time to estimate the treatment effects (see Section 3.4), our results should qualitatively hold.
@@ 3.3. Discussion Topics During COVID-19
We identify the following topics: General COVID response, COVID prevention, COVID News/statistics (general), COVID news/statistics (state-specific), Socio-Economic, Political-Religious, and China & Global. We do a qualitative and quantitative analysis of the topic clusters and find that the COVID-specific topics are very similar and merge them into a single topic COVID Response to get the final four topics. [^11] The qualitative analysis considers the most frequent words associated with each topic while quantitative analysis examines distance between cluster means and mean of pairwise cosine distances across clusters. Our results for metalearners in Section 4.2 remain the same with and without merging topic clusters. We provide the fifty most representative phrases associated with each of these topics and a sample of tweets associated with each topic in Appendix S4 Tables S3 and S4 respectively.
@@ 3.4.3. Pre-treatment Covariates
The descriptive statistics for the final event-level dataset are provided in Appendix C Table 3.
@@ 3.4.4. T-Learner
$Y(1)$ is the potential outcome for an individual if they were treated ($T=1$), i.e. if they interacted outside their group; $Y(0)$ is the potential outcome if the same individual belonged to the control group and were not treated ($T=0$); $X$ is the vector of pre-treatment covariates. We focus on $\Delta GCS$, the change in Group Conformity Score, as our main outcome. We observe either $Y(1)$ or $Y(0)$ for any given individual, and use T-Learner (Künzel et al., 2019) to estimate the unobserved potential outcome and then CATE in the following stages:
@@ 4.1. Qualitative Content Analysis based on GCS
To understand whether $GCS$ provides a more meaningful measure compared to $GCS^{BOW}$, we examine tweets by the top thirty high-$GCS$ users in both religious groups across the 7-day window post each event. All $GCS$ values in this subset lie above the $98^{th}$ percentile for each event-religion-GCS (i.e., $GCS$ or $GCS^{BOW}$) combination. We present five example tweets from these in Appendix S5 Tables S5 (high $GCS$) and S6 (high $GCS^{BOW}$). As expected, the tweets from users in one group having high group conformity are similar to tweets from users in the other group having low group conformity. This holds even without exactly matching but semantically similar tweets in the case of $GCS$. The highest $GCS^{BOW}$ tweets for the majority group often comprise a few common words. This is consistent with the results in Appendix B, Figure 4 in which we examine the relation of tweet length with $GCS^{BOW}$ and $GCS$. We find that the average tweet length is low at extreme values of $GCS^{BOW}$ while this relation is weak in the case of $GCS$.
@@ 4.2. Effect of Interaction on Change in GCS
We examine the heterogeneity in the treatment effect (TE) by regressing the treatment effect $\hat{\tau}$ on standardized pre-treatment covariates for each event. Appendix D, Figure 5 reports the complete results; we highlight the most important findings here. We find a high positive correlation between pre-treatment $GCS$ and $\hat{\tau}$. In other words, the decline in $GCS$ due to intergroup interaction is stronger for people with an already low $GCS$. This is especially true in case of the Bihar Manifesto which is a highly political event, and on which Hindus and Muslims might have divergent perspectives. However, this positive correlation breaks down for the Tablighi event. Among the topics, $\hat{\tau}$ is more negative after the launch of Coronil remedy for people who were initially more engaged in COVID response discussion, and hence, might have shared concerns related to this. The other notable topic is Politics-Religion with which $\hat{\tau}$ has a highly positive correlation in the case of Exam Satyagraha and Bihar Manifesto both of which are political events. Specifically, Exam Satyagraha was called for by the left-wing student organization AISA when the right-wing ruling government announced the decision to conduct exams. Among the emotions, we find a high positive correlation of $\hat{\tau}$ with Anger in case of the communally charged Tablighi event indicating that the intergroup interaction increased $GCS$ for people who expressed more anger earlier.
@@ 4.3. Decomposition Analysis
Given that the effects of interaction on change in $GCS$ exhibit substantial heterogeneity across the two religious groups, a natural question is—what is the contribution of topics and emotions towards explaining these differences? Importantly, emotions and topics are also computed as properties of the tweet text, and changes in $GCS$ partially embody changes in these attributes. [^15] Appendix S6, Figures S2S10 show the distribution of treatment effects on these explanatory variables across Muslims and non-Muslims for each event.
@@
Appendix D, Figure 6 decomposes the explained component into contributions of emotions and topics. For Janata Curfew, 81% of $\Delta\bar{\hat{\tau}}$ is explained by valence. This is because, for this event, $\beta^{valence}_{NM}$ is negative—i.e. an increase in valence due to intergroup interaction is associated with a decrease in $GCS$—and the difference $\bar{\hat{\tau}}^{valence}_{M}-\bar{\hat{\tau}}^{valence}_{NM}$ is positive. Additionally, 29% of $\Delta\bar{\hat{\tau}}$ is explained by the topic China & Global for this event. In contrast joy, sadness, and Politics-Religion topic have countervailing effects, i.e. they pull $\Delta\bar{\hat{\tau}}$ towards zero. In case of the Tablighi incident, valence (9%) and joy (14%) explain an important share of $\Delta\bar{\hat{\tau}}$ while anger has a countervailing contribution (-11%). For both the politically salient events—Exam Satyagraha and Bihar Manifesto—the differential effects on Politics-Religion and Socio-Economic topics explain $\Delta\bar{\hat{\tau}}$.
@@ Acknowledgements.
Appendix
Appendix A Selecting Muslim Classification Threshold
@@
Appendix B BOW-GCS and Contextualized-GCS Tweet Length
@@
Appendix C Summary Statistics
@@
Appendix D Treatment Effect Heterogeneity
@@
Appendix S2 User Name Cleaning and Filtering Organizations
We transliterate Twitter usernames from Indic languages Hindi, Bengali, Gujarati, Punjabi, Malayalam, Kannada, Tamil, Telugu, Oriya, Marathi, Assamese, Konkani, Bodo, Nepali, and Urdu to English using Indic-trans tool (Bhat et al., 2014). [^17] Available at https://github.com/libindic/indic-trans. To drop non-personal names, we construct a name part dictionary using names from a 3% random sample of eligible voters from Indian electoral rolls and the Rural Economic & Demographic Survey (REDS) data collected by the National Council of Applied Economic Research. We use the 3% sample due to data availability constraints. This is a large sample comprising over 25 million voter names and their parent’s/spouse’s names (15,431,765 unique names) out of over 800 million total voters. For every Twitter user, only name parts that occur in the constructed name part dictionary are retained. We further manually scan names of users who either have more than 20,000 followers, tweeted more than 60 times, or whose names contain any of the organization-related keywords provided in Table S2 and drop those having non-personal names from this list.
Table S2. Keywords used to filter out organization names from the tweeters after lower-casing the usernames

### 10 Mar 2024 v3

@@ Abstract.
Social media, Polarization, Intergroup Interaction, Religion
copyright: nonejournalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645675isbn: 979-8-4007-0171-9/24/05ccs: Information systems Social networksccs: Social and professional topics Religious orientationccs: Applied computing Economicsccs: Applied computing Sociology
@@ 1. Introduction
Polarization between identity groups drives social unrest and adversely affects a nation’s economic growth and responses to crises (Esteban and Schneider, 2008; Egorov et al., 2021). However, it is less clear how polarization evolves during crises. While collective suffering may foster within-group solidarity (Bauer et al., 2016), it may also lead to attribution of blame on the “outside” group or increase between-group competition for limited resources (Homer-Dixon, 2010). These behaviors might be accentuated among people who interact only within their groups and hence might have a restricted information environment (Halberstam and Knight, 2016). Polarization has been widely studied in the social-media context as well. In particular, social media platforms such as Twitter have become more polarized over time (Garimella and Weber, 2017). Consistent with this, political polarization decreased for users who deactivated their Facebook accounts before the US elections (Allcott et al., 2020). The depolarization, however, depends on an individual’s background and this pattern might reverse for individuals having homogeneous offline social networks (Asimovic et al., 2021). Interactions within social media platforms might have varying impacts on polarization. In an unfavourable environment, exposure to outgroup viewpoints may lead to stereotype formation and increase polarization. For example, Bail et al. (2018) conduct a field experiment and find that Republicans (Democrats) who were offered financial incentives to follow a liberal (conservative) Twitter bot became more conservative (liberal).
@@ 3.1. Inferring Religion
We drop verified users from our data to remove influential individuals/organizations. To reduce possible biases in the polarization measure due to erroneous classification of religion, we remove non-personal names. We discuss the details of name cleaning steps in Appendix A. Our final data comprises 692,559 unique users with 3,072,503 tweets. [^6] To check for the possibility that our results might be influenced by bots, we use the recent lists TwiBot-20 (Feng et al., 2021) and TwiBot-22 (Feng et al., 2022). We find that less than 0.5% of users in our data are listed as bots and contribute to 0.63% of the tweets. The model from (Chaturvedi and Chaturvedi, 2024) is trained on names obtained from a nationally representative sample, while Twitter usernames tend to be noisy. The name classification exercise also depends on the distribution of names in the specific domain the algorithm was trained on. Therefore, we expect a domain shift when applying the model to our data. To address this, we manually annotate a subset of approximately one thousand names as Muslim or non-Muslim. Since Muslim and non-Muslim names are linguistically distinct, prior literature has also relied on manually annotating religion as Muslim and non-Muslim from names in the South Asian context. In Appendix B, we discuss how we select the sample of names for manual annotation and choose a more suitable classification threshold. Chaturvedi and Chaturvedi (2024) report an $F_{1}$ score of 95% on their test set which reduces to around 85% when only one name part is available (i.e., only the first or last name)—as is often the case on Twitter; we find comparable values (sensitivity 84% and specificity 86.5%). The misclassifications lead to measurement error, potentially underestimating polarization; though qualitatively we expect the temporal patterns to remain the same. Therefore, our estimates are likely to be conservative estimates. Similarly, since we focus on changes in GCS over time to estimate the treatment effects (see Section 3.4), our results should qualitatively hold.
@@ 3.3. Discussion Topics During COVID-19
We identify the following topics: General COVID response, COVID prevention, COVID News/statistics (general), COVID news/statistics (state-specific), Socio-Economic, Political-Religious, and China & Global. We do a qualitative and quantitative analysis of the topic clusters and find that the COVID-specific topics are very similar and merge them into a single topic COVID Response to get the final four topics. [^11] The qualitative analysis considers the most frequent words associated with each topic while quantitative analysis examines distance between cluster means and mean of pairwise cosine distances across clusters. Our results for metalearners in Section 4.2 remain the same with and without merging topic clusters. We provide the fifty most representative phrases associated with each of these topics and a sample of tweets associated with each topic in Appendix S3 Tables S2 and S3, respectively.
@@ 3.4.3. Pre-treatment Covariates
The descriptive statistics for the final event-level dataset are provided in Appendix D Table 4.
@@ 3.4.4. T-Learner
Where $Y(1)$ is the potential outcome for an individual if they were treated ($T=1$), i.e. if they interacted outside their group; $Y(0)$ is the potential outcome if the same individual belonged to the control group and were not treated ($T=0$); $X$ is the vector of pre-treatment covariates. We focus on $\Delta GCS$, the change in Group Conformity Score, as our main outcome. We observe either $Y(1)$ or $Y(0)$ for any given individual, and use T-Learner (Künzel et al., 2019) to estimate the unobserved potential outcome and then CATE in the following stages:
@@ 4.1. Qualitative Content Analysis based on GCS
To understand whether $GCS$ provides a more meaningful measure compared to $GCS^{BOW}$, we examine tweets by the top thirty high-$GCS$ users in both religious groups across the 7-day window post each event. All $GCS$ values in this subset lie above the $98^{th}$ percentile for each event-religion-GCS (i.e., $GCS$ or $GCS^{BOW}$) combination. We present five example tweets from these in Appendix S4 Tables S4 (high $GCS$) and S5 (high $GCS^{BOW}$). As expected, the tweets from users in one group having high group conformity are similar to tweets from users in the other group having low group conformity. This holds even without exactly matching but semantically similar tweets in the case of $GCS$. The highest $GCS^{BOW}$ tweets for the majority group often comprise a few common words. This is consistent with the results in Appendix C, Figure 4 in which we examine the relation of tweet length with $GCS^{BOW}$ and $GCS$. We find that the average tweet length is low at extreme values of $GCS^{BOW}$ while this relation is weak in the case of $GCS$.
@@ 4.2. Effect of Interaction on Change in GCS
We examine the heterogeneity in the treatment effect (TE) by regressing the treatment effect $\hat{\tau}$ on standardized pre-treatment covariates for each event. Appendix E, Figure 5 reports the complete results; we highlight the most important findings here. We find a high positive correlation between pre-treatment $GCS$ and $\hat{\tau}$. In other words, the decline in $GCS$ due to intergroup interaction is stronger for people with an already low $GCS$. This is especially true in case of the Bihar Manifesto which is a highly political event, and on which Hindus and Muslims might have divergent perspectives. However, this positive correlation breaks down for the Tablighi event. Among the topics, $\hat{\tau}$ is more negative after the launch of Coronil remedy for people who were initially more engaged in COVID response discussion, and hence, might have shared concerns related to this. The other notable topic is Politics-Religion with which $\hat{\tau}$ has a highly positive correlation in the case of Exam Satyagraha and Bihar Manifesto both of which are political events. Specifically, Exam Satyagraha was called for by the left-wing student organization AISA when the right-wing ruling government announced the decision to conduct exams. Among the emotions, we find a high positive correlation of $\hat{\tau}$ with Anger in case of the communally charged Tablighi event indicating that the intergroup interaction increased $GCS$ for people who expressed more anger earlier.
@@ 4.3. Decomposition Analysis
Given that the effects of interaction on change in $GCS$ exhibit substantial heterogeneity across the two religious groups, a natural question is—what is the contribution of topics and emotions towards explaining these differences? Importantly, emotions and topics are also computed as properties of the tweet text, and changes in $GCS$ partially embody changes in these attributes. [^15] Appendix S5, Figures S2–S10 show the distribution of treatment effects on these explanatory variables across Muslims and non-Muslims for each event.
@@
Appendix E, Figure 6 decomposes the explained component into contributions of emotions and topics. For Janata Curfew, 81% of $\Delta\bar{\hat{\tau}}$ is explained by valence. This is because, for this event, $\beta^{valence}_{NM}$ is negative—i.e. an increase in valence due to intergroup interaction is associated with a decrease in $GCS$—and the difference $\bar{\hat{\tau}}^{valence}_{M}-\bar{\hat{\tau}}^{valence}_{NM}$ is positive. Additionally, 29% of $\Delta\bar{\hat{\tau}}$ is explained by the topic China & Global for this event. In contrast joy, sadness, and Politics-Religion topic have countervailing effects, i.e. they pull $\Delta\bar{\hat{\tau}}$ towards zero. In case of the Tablighi incident, valence (9%) and joy (14%) explain an important share of $\Delta\bar{\hat{\tau}}$ while anger has a countervailing contribution (-11%). For both the politically salient events—Exam Satyagraha and Bihar Manifesto—the differential effects on Politics-Religion and Socio-Economic topics explain $\Delta\bar{\hat{\tau}}$.
@@ Acknowledgements.
Appendix A User Name Cleaning and Filtering Organizations
We transliterate Twitter usernames from Indic languages Hindi, Bengali, Gujarati, Punjabi, Malayalam, Kannada, Tamil, Telugu, Oriya, Marathi, Assamese, Konkani, Bodo, Nepali, and Urdu to English using Indic-trans tool (Bhat et al., 2014). [^16] Available at https://github.com/libindic/indic-trans. To drop non-personal names, we construct a name part dictionary using names from a 3% random sample of eligible voters from Indian electoral rolls and the Rural Economic & Demographic Survey (REDS) data collected by the National Council of Applied Economic Research. We use the 3% sample due to data availability constraints. This is a large sample comprising over 25 million voter names and their parent’s/spouse’s names (15,431,765 unique names) out of over 800 million total voters. For every Twitter user, only name parts that occur in the constructed name part dictionary are retained. We further manually scan names of users who either have more than 20,000 followers, tweeted more than 60 times, or whose names contain any of the organization-related keywords provided in Table 3 and drop those having non-personal names from this list.
Table 3. Keywords used to filter out organization names from the tweeters after lower-casing the usernames
Appendix B Muslim Classification Threshold
@@
Appendix C Tweet Length: BOW-GCS and GCS
@@
Appendix D Summary Statistics
@@
Appendix E Treatment Effect Heterogeneity